# Driving Scene Synthesis on Free-form Trajectories with Generative Prior

### v1

@@ Abstract
Driving scene synthesis along free-form trajectories is essential for driving simulations to enable closed-loop evaluation of end-to-end driving policies. While existing methods excel at novel view synthesis on recorded trajectories, they face challenges with novel trajectories due to limited views of driving videos and the vastness of driving environments. To tackle this challenge, we propose a novel free-form driving view synthesis approach, dubbed DriveX, by leveraging video generative prior to optimize a 3D model across a variety of trajectories. Concretely, we crafted an inverse problem that enables a video diffusion model to be utilized as a prior for many-trajectory optimization of a parametric 3D model (e.g., Gaussian splatting). To seamlessly use the generative prior, we iteratively conduct this process during optimization. Our resulting model can produce high-fidelity virtual driving environments outside the recorded trajectory, enabling free-form trajectory driving simulation. Beyond real driving scenes, DriveX can also be utilized to simulate virtual driving worlds from AI-generated videos.
Figure 1: Though conventional reconstruction method can well fit views on the recorded trajectory, they struggle to extrapolate onto the views along the novel trajectories (top). In contrast, our method significantly improves the synthesis quality on these novel views (bottom), thereby achieving driving scene synthesis on free-form trajectories.
@@ 1 Introduction
Building virtual driving worlds engagible with driving policies plays a pivotal role in developing robust autonomous driving systems. It enables automatic synthesis of diverse driving data, including safety-critical long-tail scenarios, edge cases, and providing closed-loop evaluation for end-to-end autonomous driving systems.
Recent advancements in novel view synthesis [21, 15] have driven reconstruction-based driving simulation [34, 4, 35] due to its ability to synthesize photorealistic sensor data. These methods aim to reconstruct driving environments from individual driving sequences, typically involving single-trajectory videos from surrounding cameras with sparse views, minimal overlap, and textureless areas. Despite impressive results in fitting training views and interpolating along original trajectories, they still struggle to extrapolate substantially novel views far from the recorded trajectory, leading to limited flexibility and usefulness for driving simulation. To address these problems, a couple of latest works [13, 38] explored image generative prior as extra novel view regularization. However, they are largely limited in design due to (i) the need for fine-tuning an off-the-shelf image generation models to obtain the capability of view transformation, suffering from both additional computational cost and per-scene data scarcity, and (ii) absence of the desired spatiotemporal knowledge within image models. To tackle these issues, a concurrent work [40] uses a pretrained video diffusion model to generate novel trajectory training images conditioned on initial recorded frames. Without sufficient content constraint but with intrinsic hallucination risk, this approach however tends to drift away from the original scene in scene generation, making troubles for the subsequent model optimization.
To address the challenges mentioned above, we propose a novel driving scene synthesis approach for free-form trajectories, termed DriveX, by leveraging video generative prior with rich spatiotemporal knowledge during the model (e.g., Gaussian splatting) optimization. This is achieved by designing an inverse problem that enables a video diffusion model to serve as a prior for optimizing the Gaussian over a variety of trajectories. During training, we render novel trajectory views using the in-training model. To identify potential artifacts from trajectory extrapolation, we compare each rendered view with recorded images wrapped under the same view and geometric conditions, allowing us to obtain reliable regions for each rendered image. The inverse problem is then crafted as restoring each rendered view via utilizing its corresponding reliable regions as conditioning inputs, which can be effectively executed by the video diffusion model. The output from this process is subsequently fed into the optimization as supervision to update the target model. To seamlessly exploit the generative prior, we iteratively conduct this process during Gaussian optimization. As shown in Fig. 1, our approach substantially improve the view synthesis quality in novel trajectory compared to the traditional reconstruction methods.
The contributions of this paper are as follows: (i) We propose leveraging video generative prior with rich spatiotemporal knowledge for generalizable driving scene synthesis from single-trajectory recorded videos. (ii) We introduce DriveX, a novel driving scene synthesis framework for free-form trajectories that innovatively constructs an inverse problem to enable the use of a video diffusion model as a prior. (iii) Extensive experiments demonstrate that DriveX significantly outperforms existing state-of-the-art alternatives in driving scene synthesis using single-trajectory recorded videos. Additionally, we showcase the superiority of our method in rendering novel trajectories of virtual driving worlds from AI-generated videos, even when faced with inherent content inconsistencies. This approach enables a more economically scalable large-scale simulation without the need to collect videos for each scene.
@@ 2 Related works
Early approaches [34, 36, 27, 32, 28] for dynamic urban scenes reconstruction focus on extending neural radiance fields (NeRF) [21] in large scale unbounded scenes.These methods are restricted with low efficiency and blurry rendering performance. Recently, another line of works introduced 3D Gaussian splatting [15] into this task to produce real-time high-fidelity rendering. Some of them [35, 42] leverage its favorable explicit property to decouple the whole scene into the static background and moving foreground Gaussians, while others modeling the motion of 3D Gaussians with periodic functions [4] or deformation fields [12]. However, driving scenes featured with sparse views, little overlaps, and large textureless areas, making existing methods struggle to generalize to substantively novel views.
Generative-based driving scene synthesis
@@
Recently, propelled by the significant advancement of the diffusion model [10, 24, 22, 25, 1], many works employ video generators to construct world models [16]. Some of them introduce this paradigm into autonomous driving [29, 41, 17, 6, 30, 7, 11, 20] for simulating the plausible visual outcomes under versatile driving controls. Despite the remarkable diversity of generated results, this approach is constrained by the absence of underlying 3D model. Consequently, they cannot promise geometry and texture consistency for the generated many-trajectory videos of the same scene.
@@
Figure 2: The schematic illustration of DriveX. Top: Gaussian optimization with both recorded trajectory and generated novel trajectory as supervision. Bottom: Generative prior integration. Given a novel view rendered image from Gaussians $\mathcal{G}_{t}$, video diffusion is applied to generate its refined version with condition of unreliability mask. We omit the loss on recorded trajectory for clarity.
Scene reconstruction with diffusion prior
Leveraging generative prior to enhance the reconstruction from sparse view observations [37, 19, 18] has emerged. Extending this paradigm into driving scene reconstruction has been recently attempted [38, 13, 40]. However, they typically suffer from limited diversity and scale of driving scene data. SGD [38] and VEGS [13] contribute to define and tackle pseudo and extrapolated view synthesis in driving scenarios. However, these methods may encounter challenges with accurate pose and detail estimation under novel trajectories. GGS [8] introduces virtual lane generation module, but its dependency on modified MVSplat structure poses challenges for generalization to large-scale pretraining. A concurrent work [40] uses a world model [29] to generate novel trajectory video referenced on initial frame for trajectory augmentation. However, it tends to drift away from the original scene due to the strong built-in hallucination risk and insufficient constraints. This would inevitably hurt the subsequent model optimization. In contrast, our method dedicates to solve an inverse problem that enables a video diffusion model to be used as a prior for many-trajectory optimization of a parametric 3D model, achieving superior synthesis quality on free-form trajectories.
@@ 3.1 Preliminary: Driving scene reconstruction
Reconstruction-based driving simulation aims to recover the surrounding scenes from calibrated video $\mathcal{V}_{\mathrm{gt}}=\left\{I_{i}\right\}_{i=0}^{F}$ recorded along a driving trajectory $\mathcal{T}=\left\{P_{i}\right\}_{i=0}^{F}$, where $P_{i}\in\mathrm{SE(3)}$ is the camera pose at frame $i$. Then the sensor data would be rendered when the vehicle performs simulated unrecorded actions (e.g. lane change) in the reconstructed environment.
@@
In the urban scenes, although the traffic participants are dynamic, they can be regarded as static in their own local coordinate frames. The transformation from each object’s local frame to the global coordinate frame of the whole scene is parameterized by a series of rotations and transformations, which can be derived from its tracklet annotated by off-the-shelf 3D trackers [5, 33]. Then a complete dynamic street scene can be modeled by decoupled foreground moving Gaussians and static background Gaussians. Additionally, cube map is commonly used to model the sky with infinite distance [35, 4].
Challenges
Since the camera configuration in autonomous vehicles is primarily designed for perception tasks, the captured videos often offer limited viewpoints, minimal overlap, and large textureless regions. These inherent properties bring challenges to the reconstruction of the unbounded driving scenes from single-trajectory videos. Fitting the training views at recorded trajectory $\mathcal{T}$ is insufficient to realize satisfactory extrapolation out of $\mathcal{T}$ for novel view synthesis. As a result, existing optimization-based driving scene synthesis approaches struggle to achieve high-quality view rendering along novel trajectories $\mathcal{T}^{\prime}=\{P^{\prime}_{i}|P^{\prime}_{i}\notin\mathcal{T}\}_{i=0}^{F}$, often leading to obvious degeneration and artifacts.
3.2 DriveX
To address the aforementioned challenges, we propose a novel framework, DriveX, seamlessly integrating generative priors into the driving scene reconstruction process (e.g. Gaussian optimization). As shown in Fig. 2, it iteratively synthesizes novel trajectory scenes by utilizing a video diffusion as supervision to optimize the Gaussian model. To ensure the generated content aligns with the underlying scenes in detail, the generation process is appropriately conditioned on the in-training Gaussian model.
To achieve this, we delicately design an inverse problem [14, 23], which can be formulated as recovering clean novel view images $\mathcal{V}$ from the artifact-heavy novel view rendered images $\mathcal{V}^{\prime}$ on novel trajectories $\mathcal{T}^{\prime}$. Concretely, $\mathcal{V}^{\prime}$ can be regarded as a measurement of unknown ground truth $\mathcal{V}$, as:
$\mathcal{V}^{\prime}=f(\mathcal{V})+\epsilon,$ (3)
where $\epsilon$ is a random noise, and the measurement function $f$ distorts $\mathcal{V}$ due to poor generalization ability of Gaussian model in conventional methods. With this design, a video generative model [37] can be readily employed to address this problem effectively. The refined video can then serve as additional supervision within Gaussian optimization.
Solving inverse problem with generative prior
The key to solve the inverse problem lies in accurately identifying the artifacts in the novel view rendered video. Inspired from the sparse view reconstruction methods [34, 19], we construct an unreliability mask representing the artifacts via comparing the rendered images with recorded images, which is used to directly guide the diffusion model focus on refining these unreliable regions while preserving reliable areas.
Specifically, given a sequence of viewpoints $\mathcal{T}^{\prime}=\left\{P_{i}^{\prime}\right\}_{i=0}^{F}$ along a novel trajectory, a video diffusion model $\mathcal{D}$ is employed to estimate $\mathcal{V}_{i}$ in Eq. 3:
@@ 3.2 DriveX
where $\mathcal{V}^{\prime}_{t}$ is rendered from the Gaussian model $\mathcal{G}_{t}$ at current iteration $t$ through Eq. 1, the refined video $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ can be regard the estimation of $\mathcal{V}$ in Eq. 3, and $\mathcal{M}$ is the mask indicating unreliable regions. Typically, $\mathcal{D}$ takes noise-perturbed images as input, conditioned on the masked video, and outputs the refined video $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$. Then $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ is used as the additional supervision for the novel trajectory loss $\mathcal{L}^{\prime}$ (Eq. 6) to improve the novel view rendering. Note we omit the loss on recorded trajectory [35] in the whole context for clarity.
@@
To assess geometry accuracy in novel view rendered images, we derive $\mathcal{M}$ by comparing the rendered images with novel view warped images from the recorded trajectory. Let $I_{\text{ren}}\in\mathcal{V}^{\prime}_{t}$ denote a rendered image, and let $I_{\text{rec}}$ be the recorded image that are closest to $I_{\text{ren}}$. We define a warping operation $\psi$ in Eq. 7 to project a 3D point $p$ onto the image plane given camera pose $P$:

### v2

# DriveX: Driving View Synthesis on Free-form Trajectories with Generative Prior

@@ Abstract
Driving view synthesis along free-form trajectories is essential for realistic driving simulations, enabling closed-loop evaluation of end-to-end driving policies. Existing methods excel at view interpolation along recorded paths but struggle to generalize to novel trajectories due to limited viewpoints in driving videos. To tackle this challenge, we propose DriveX, a novel free-form driving view synthesis framework, that progressively distills generative prior into the 3D Gaussian model during its optimization. Within this framework, we utilize a video diffusion model to refine the degraded novel trajectory renderings from the in-training Gaussian model, while the restored videos in turn serve as additional supervision for optimizing the 3D Gaussian. Concretely, we craft an inpainting-based video restoration task, which can disentangle the identification of degraded regions from the generative capability of the diffusion model and remove the need of simulating specific degraded pattern in the training of the diffusion model. To further enhance the consistency and fidelity of generated contents, the pseudo ground truth is progressively updated with gradually improved novel trajectory rendering, allowing both components to co-adapt and reinforce each other while minimizing the disruption on the optimization. By tightly integrating 3D scene representation with generative prior, DriveX achieves high-quality view synthesis beyond recorded trajectories in real time—unlocking new possibilities for flexible and realistic driving simulations on free-form trajectories.
Figure 1: Though conventional reconstruction-based approaches can well fit views on the recorded trajectory, they struggle to extrapolate onto the views beyond it (top). In contrast, our method significantly improves the synthesis quality on these novel trajectory views by incorporating generative prior (bottom), thereby achieving superior synthesis quality on free-form trajectories.
@@ 1 Introduction
Building virtual driving worlds capable of being engaged with driving policies plays a pivotal role in developing robust autonomous driving systems. It facilitates efficient training data scaling and automatic synthesis of safety-critical long-tail cases, while enabling close-loop evaluation for end-to-end autonomous driving systems.
In recent years, the burgeoning advancements in novel view synthesis [19, 13] have propelled the prevalence of reconstruction-based driving simulations [36, 3, 37]. These methods aim to reconstruct driving environments from vehicle sensor data typically involving single-trajectory videos from surrounding cameras and synchronized LiDAR point clouds. However, the limited overlap between frames, large textureless regions and illumination variations in large-scale scenes, drastically complicate this task than conventional novel view synthesis. Consequently, despite achieving impressive results in fitting training views and interpolating along recorded trajectories, they still struggle to extrapolate on novel views far from the recorded trajectory, limiting their flexibility and usefulness in many downstream tasks.
Recent works [16, 17, 39] have demonstrated that 2D generative priors can effectively regularize sparse-view reconstruction by synthesizing pseudo ground-truth images for novel view, thereby alleviating the under-constraint nature of this task. However, directly applying these approaches in the driving scenes presents unique challenges due to the limited diversity of trajectories in existing driving video datasets. The training is primarily confined on forward-moving trajectories, while the off-trajectory videos are required for inference. This will introduce a considerable gap, making it infeasible to learn precise explicit pose control, and posing a great challenge to simulate degraded patterns for training a restoration-based model.
To address these challenges, we propose a reconstruction-generation interwined framework for driving view synthesis along free-form trajectories, termed DriveX, which progressively distills video generative prior into the 3D scene representation (i.e., 3D Gaussian Splatting [13]). During this process, a video diffusion model is employed to generate novel trajectory videos as additional supervision for the optimization of scene representation, effectively mitigating the sparse-view issue. This is achieved by leveraging a generative model to restore novel trajectory renderings from the in-training 3D Gaussian model, which ensures the consistency of generated content while eliminating the need for explicit pose condition. Specifically, we design an inpainting-based restoration task where generation is conditioned on the rendered video masked by a geometry-aware unreliability mask. This formulation allows the model to fully leverage its generative capability, freeing it from the task of identifying degraded regions, thus circumventing the complexity of simulating degraded patterns during its training. During the optimization of scene representation, we progressively perform the restoration process and update the generated pseudo ground truth to further enhance consistency of the generated content. As shown in Fig. 1, our approach substantially improves the view synthesis quality in the novel trajectory compared to the state-of-the-art reconstruction-only method [37].
Our contributions are as follows: (i) We propose a novel driving view synthesis framework that seamlessly integrates generative prior into driving scene reconstruction, where a diffusion model is employed to progressively synthesize consistent novel view ground truth for regularizing the optimization of the 3D Gaussian model; (ii) By formulating an inpainting-based restoration task, we disentangle the identification of degraded regions from the generative capability of the diffusion model, thereby simplifying training on driving videos with limited diverse trajectories; (iii) Our DriveX effectively overcomes the sparse-view issue in previous street scene reconstruction methods, enabling high-fidelity view synthesis along free-form trajectories.
@@ 2 Related works
Reconstruction-based driving view synthesis   Early approaches [36, 38, 26, 32, 27] for dynamic urban scenes reconstruction focus on extending neural radiance fields (NeRF) [19] in large scale unbounded scenes. These methods are restricted with low efficiency and blurry rendering performance. Recently, another line of works introduced 3D Gaussian splatting [13] into this task to produce real-time high-fidelity rendering. Some of them [37, 46] leverage its favorable explicit property to decouple the whole scene into the static background and moving foreground Gaussians, while others modeling the motion of 3D Gaussians with periodic functions [3] or deformation fields [11]. However, driving scenes featured with sparse views, little overlaps, and large textureless areas, making existing methods struggle to generalize to substantively novel views.
@@
Generative-based driving view synthesis   Recently, propelled by the significant advancement of the diffusion model [9, 23, 22, 24, 1], many works employ video generators to construct world models [14]. Some of them introduce this paradigm into autonomous driving [29, 45, 15, 5, 30, 6, 10, 18] for simulating the plausible visual outcomes under versatile driving controls. Despite the remarkable diversity of generated results, this approach is constrained by the absence of underlying 3D model. Consequently, they cannot promise geometry and texture consistency for the generated many-trajectory videos of the same scene.
@@
Figure 2: Schematic illustration of DriveX. Top: Optimizing the scene representation with both recorded trajectory views and generated novel trajectory views as supervision. Bottom: Generative prior integration. During optimization, given a novel view rendering from in-training Gaussian $\mathcal{G}_{t}$, a video diffusion model is employed to generate a restored version as pseudo ground truth on novel trajectories, guided by its unreliability mask and accumulated LiDAR projections.
Scene reconstruction with diffusion prior   Leveraging generative prior to enhance the reconstruction from sparse view observations [39, 17, 16, 34] has emerged. Extending this paradigm into driving scene reconstruction has been recently attempted [40, 12, 44]. Limited diversity of trajectory in driving scene data poses unique challanges for them. SGD [40] and VEGS [12] contribute to define and tackle pseudo and extrapolated view synthesis in driving scenarios. However, these methods may encounter challenges with accurate pose and detail estimation under novel trajectories. DriveDreamer4D [44] samples novel trajectory video based on single reference frame before training, risking hallucination and inconsistency due to insufficient constraints. StreetCrafter [41] refines the novel view renderings using a controllable video diffusion model guided by LiDAR projections. Meanwhile, FreeSim [4] and Difix3D+ [33] directly condition on degraded renderings and employ carefully data curation to train the generative model to adapt on degradation patterns resulted from off-trajectory renderings. In contrast, by crafting an inpainting-based video restoration task, we free the generative model from identifying degraded regions, thereby eliminating the need to simulate specific degraded patterns in training.
@@ 3.1 Preliminary: Driving scene reconstruction
Reconstruction-based driving simulation aims to recover the surrounding scenes from calibrated video $\mathcal{V}_{\mathrm{gt}}=\left\{I_{i}\right\}_{i=0}^{F}$ recorded along a driving trajectory $\mathcal{T}=\left\{P_{i}\right\}_{i=0}^{F}$, where $P_{i}\in\mathrm{SE(3)}$ is the camera pose at frame $i$. Then the sensor data can be simulated under the unrecorded actions (e.g. lane change) in the reconstructed environment.
@@
Following [37], we model the traffic participants in the driving scenes as a dynamic node in the neural scene graph with rigid motion, which can be regarded as static in their own local coordinate frames. The transformation from each object’s local frame to the global coordinate frame of the whole scene is parameterized by a series of rotations and transformations, which can be annotated by their tracked bounding boxes. Additionally, cube map is employed to model the sky with infinite distance [37, 3].
Challenges   Since the camera configuration in autonomous vehicles is primarily designed for perception tasks, the captured videos often offer limited viewpoints, minimal overlap, and large textureless regions. These inherent properties bring challenges to the reconstruction of the unbounded driving scenes from single-trajectory videos. Fitting the training views at recorded trajectory $\mathcal{T}$ is insufficient to realize satisfactory extrapolation out of $\mathcal{T}$ for novel view synthesis. As a result, existing optimization-based driving scene synthesis approaches struggle to achieve high-quality rendering along novel trajectories $\mathcal{T}^{\prime}=\{P^{\prime}_{i}|P^{\prime}_{i}\notin\mathcal{T}\}_{i=0}^{F}$, often leading to obvious degeneration and artifacts.
3.2 Novel view regularization with generative prior
Generate consistent novel view via video restoration A promising solution to the aforementioned challenges is to incorporate generative priors as an additional regularization. Large video generative models trained on internet-scale data have developed a strong knowledge of natural video distributions, enabling them to reliably infer and generate novel view content as pseudo ground truth for optimizing 3DGS. To ensure consistency between the generated content and the underlying scene, thereby preventing disruption to optimization, the generation should be appropriately conditioned on the in-training Gaussian model.
To achieve this, we delicately design a video restoration task, which is formulated as recovering clean novel view images $\mathcal{V}$ from the artifact-heavy novel view renderings $\mathcal{V}^{\prime}$ beyond the recorded trajectory. Within this design, a video generative model can be employed to address this problem effectively. The refined video then serves as additional supervision in the optimization of 3DGS.
Concretely, given a sequence of viewpoints $\mathcal{T}^{\prime}=\left\{P_{i}^{\prime}\right\}_{i=0}^{F}$ along a novel trajectory, a video diffusion model $\mathcal{D}$ is employed to estimate clean version $\mathcal{V}_{i}$:
@@ 3.2 Novel view regularization with generative prior
$\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ $=\mathcal{D}\left(\mathcal{V}^{\prime}_{t}\right),$ (4)
@@
where $\mathcal{V}^{\prime}_{t}$ is rendered from the Gaussian model $\mathcal{G}_{t}$ at $t$-th iteration of optimization through Eq. 1, the refined video $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ can be regarded as the estimation of $\mathcal{V}$. Typically, $\mathcal{D}$ takes noise-perturbed images as input, conditioned on the degraded video, and outputs the refined video $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$. Then $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ is used as the pseudo ground truth in the novel trajectory loss $\mathcal{L}^{\prime}$ (Eq. 5) to supervise the novel view renderings as illustrated in Fig. 2.
@@
Iterative refinement The generative model can be employed to provide novel view supervision for enhancing the novel view renderings. In turn, the generation quality also benefits from these improvements, as it is conditioned on $\mathcal{G}_{t}$ whose enhancements could provide more accurate information to guide the generation process. This co-adapt and reinforced enhancement motivates an iterative refinement in our framework. As outlined in Algorithm 1, we initiate the optimization for $T_{0}$ steps using conventional reconstruction-based method [37]. Subsequently, we perform video restoration in Eq. 4 based on the in-training Gaussians $\mathcal{G}_{T_{0}}$, and use a buffer to store refined videos for reuse in the following $K$ iterations. To ensure that the generated content can be improved along with the progressively optimized reconstruction results, we update all videos in the buffer every $K$ iterations by performing video restoration conditioned on the latest novel trajectory renderings.
Due to the additional time demands of running the video diffusion model, it is inefficient to set $K=1$, i.e., the buffer is updated in every iteration. We set it to $3000$ to strike the balance between quality and efficiency. Please refer to Sec. 4.4 for the trade-off between different choice of $K$. As the generated contents are conditioned on Gaussian model $\mathcal{G}_{t}$, they are constrained with the recorded scene, thus ensuring the consistency between the recorded and generated ground truth and leading to stable optimization.
Novel trajectory sampling Another important design space is how to sample the novel trajectory $\mathcal{T}^{\prime}=\left\{P_{i}^{\prime}\right\}_{i=0}^{F}$ to be distilled by video diffusion model. The ideal trajectory should balance the two requirements: (i) maximizing the utility of the generative model, i.e., the generated novel views should contribute as much as possible to enhance reconstruction quality; (ii) minimizing inconsistency between the recorded and generated images to mitigate potential perturbation for Gausssian’s optimization. Therefore, we adopt a panning camera trajectory that starts from a recorded front view $P^{\prime}_{0}=[R_{0}|T_{0}]$ and gradually shifts laterally:
$P^{\prime}_{i}=[R_{0}|\frac{i}{F}s\mathbf{v}+T_{0}],$ (6)
where $\mathbf{v}\in\mathbb{R}^{3}$ is the shifting direction, and $s\in\mathbb{R}$ controls the maximum shifting length. This design not only provides novel views beyond the recorded trajectory, but also allows video generation model to extract detailed references from the initial frame. At the early stage of optimization, rendering quality may degrade as the camera deviates from the recorded driving trajectory, which offers limited guidance for the video generation. To address this issue, we initially use a small shifting length $s$ and progressively extend it during optimization, ensuring our method’s stability and robustness across diverse scenes.
3.3 Inpainting-based video restoration
Under the proposed trajectory sampling scheme with progressively increasing shifted length, degradations in the rendered video caused by the poor extrapolation capabilities of the scene representation are typically localized. Therefore, the key to effective video restoration lies in accurately identifying the artifacts in the rendered videos. Ideally, a well-trained generative model should be capable of automatically identifying and restoring these degraded regions. However, limited viewpoints in driving video datasets make it challenging to simulate degradation caused by off-trajectory rendering while obtaining corresponding ground truth during the training of the generative model. Therefore, we propose an inpainting-based restoration framework that removes the need for the model to adapt to specific degraded patterns, thereby simplifying the training process.
Specifically, we construct an unreliability mask representing the artifacts via comparing the rendered images with the recorded images, which directly guides the model to focus on refining these unreliable regions while preserving reliable areas. Thus, the Eq. 4 should be revised as:
where $\mathcal{M}$ is the mask indicating unreliable regions, and $\mathcal{D}$ is conditioned on the masked video.
Unreliability mask To assess geometry accuracy in novel view rendered images, we derive $\mathcal{M}$ by comparing the rendered images with novel view warped images from the recorded trajectory. Let $I_{\text{ren}}\in\mathcal{V}^{\prime}_{t}$ denote a rendered image, and let $I_{\text{rec}}$ represent the recorded image that are “closest” to $I_{\text{ren}}$. We define a warping operation $\psi$ in Eq. 8 to project a 3D point $p$ onto the image plane given camera pose $P$: